A ball is drawn at random from a box containing 5 blue balls, 3 green balls, and 2 black balls. Determine the probability that the ball drawn is black.
1
Expert's answer
2020-05-11T18:35:21-0400

P=n/m, where n is the number of black balls, m is the total amount of balls,

P=2/(5+3+2)=0.2